# Evaluating crop rotations around the world using satellite imagery and causal machine learning

### v1

@@ Abstract
Building sustainable food systems that are resilient to climate change will require improved agricultural management and policy. One common practice that is well-known to benefit crop yields is crop rotation, yet there remains limited understanding of how the benefits of crop rotation vary for different crop sequences and under different weather conditions. To address these gaps, we leverage crop type maps, satellite data, and causal machine learning to study how precrop effects on subsequent yields vary with cropping sequence choice and weather. Complementing and going beyond what is known from randomized field trials, we find that (i) for those farmers who do rotate, the most common precrop choices tend to be among the most beneficial, (ii) the effects of switching from a simple rotation (which alternates between two crops) to a more diverse rotation were typically small and sometimes even negative, (iii) precrop effects tended to be greater under rainier conditions, (iv) precrop effects were greater under warmer conditions for soybean yields but not for other crops, and (v) legume precrops conferred smaller benefits under warmer conditions. Our results and the methods we use can enable farmers and policy makers to identify which rotations will be most effective at improving crop yields in a changing climate.
@@ 1 Introduction
Crop rotation is a well-established agronomic practice that typically raises crop yields while reducing the need for fertilizer and pesticide inputs. Yield benefits can arise from several mechanisms, including suppression of harmful weeds, pests, and diseases [1, 2], increase of soil availability of essential nutrients such as nitrogen [3, 4, 5], and improved soil health and microbial biomass [6, 7]. Although these mechanisms are well established, less is known about how the benefits of crop rotation depend on the specific cropping sequence and on the weather conditions. This knowledge gap makes it difficult to assess how sub-optimal current farmer practices are (in many regions farmers often grow the same crop each year or use a simple 2-crop rotation) and whether shifts in crop rotation could help farmers adapt to ongoing climate changes. Answering these questions is crucial but will require moving beyond traditional research approaches that leverage data from multiple randomized field experiments.
@@
Recent advances in crop type mapping, satellite imagery, and causal machine learning have enabled researchers to assess the benefit of crop rotations using much larger sample sizes which allow examination of how rotation benefits vary with weather, rotational diversity, and the specific precrop and outcome crop. A number of recent studies used crop type maps and satellite imagery to assess the benefit of crop rotations [12, 13, 14, 15, 16, 17, 18, 19, 20], and the accuracy of some of these observational approaches have been validated against data from randomized field experiments [12] or expert recommendations [18]. However, all of these studies have limitations in terms of their scope, their methodology, or both (see Table A5 for details). In summary, none of these studies assess the impact of switching from a simple rotation to a diverse rotation, and only a few of them assess how the benefit of rotation varies with weather [12, 19, 17, 20]. Moreover, most of these studies do not use a formal causal inference framework, and instead often rely on simple comparisons between rotated and non-rotated samples. Only [12, 19, 20] use a formal causal inference framework by leveraging modern causal machine learning tools. Of these studies, only [20] quantifies the uncertainty in their estimates of how the benefit of rotation varies with weather. As with the current study, most of these studies focus on precrop effects (defined as the effect of the previous year’s crop type on the yield of the current crop) rather than total or long-term yield effects of rotation on all crops in the sequence. The former is a major (but not the only) component of rotation effects, while the latter are difficult to detect given the limited number of years in which large-scale, high-resolution crop type maps are available.
@@ 2.1 Estimates of impacts of 2-year crop sequences
In each of four countries, for each common combination of precrop A and outcome crop B, we estimated the impact of crop sequences where A is grown in the first year and B in the second year (denoted by AB) compared to no rotation controls where B was grown both years (denoted by BB). In particular, we estimated the impact of such cropping sequences on a satellite-based greenness proxy for the yield of B in the second year using a causal machine learning algorithm (see Sections 4.1-4.4 for details), and we subsequently calibrated these effect estimates to be expressed on the crop yield scale (Section 4.7).
The results (Table A8 (Column 6); Figure 3) indicate that rotations generally benefit crop yields, but there are considerable differences in these benefits for different precrop and outcome crop combinations. Of the four outcome crops considered, the rotation benefit estimates were generally largest for spring wheat and smallest for corn. For winter wheat outcome crops, soybean precrops had a negative estimated effect whereas rapeseed or fallow precrops had a positive estimated effect. For spring wheat outcome crops, legume precrops and rapeseed precrops were generally the most effective, whereas for soybean outcome crops, cereals were the most effective types of precrops (Figure A1).
The results also suggest that beneficial precrops are being exploited by farmers. In Figure 3, we see that the most common precrops have either the highest or nearly the highest benefits. There are some exceptions in which less common precrop choices are estimated to be more effective than the most common precrop choices when spring wheat or corn is the outcome crop. In addition, only relatively common precrops are included in the analysis (Section 4.3), so our analysis does not rule out the possibility that some very rare precrops are more effective than commonly used precrops.
@@ 2.2 Variability of rotation benefits with weather
We used a causal machine learning approach (causal forests) to estimate the impacts of each 2-year rotation as a nonlinear function of covariates such as weather, time, and geolocation, and we subsequently fit a linear regression of these rotation impacts on time, growing season temperature, and growing season precipitation. We then rescaled the regression coefficients for temperature and precipitation, so that they can roughly be interpreted as giving estimates of the change in precrop effect on subsequent crop yield when the growing season temperature or precipitation increases from its 25th percentile to its 75th percentile (see Sections 4.5 and 4.7 for details).
The results (Table A8 (Column 8–9); Figure 4) indicate that at higher growing season precipitation there is generally a greater benefit due to rotation. In particular, with a few exceptions in France and for fallow precrops, cases where the rotation benefit is greater at lower precipitation levels did not have statistically significant heterogeneity coefficients. In addition, based on a sensitivity analysis where we use the log of estimated crop yield as the outcome variable (Appendix A.3; Figure A4, bottom left panel), we find that this result is not merely driven by the fact that crop productivity is higher in rainier conditions. Instead, we see increases in both the relative effectiveness and absolute effectiveness of crop rotation in rainier conditions. For the rotations with a legume or broadleaf precrop, the heterogeneity of the rotation benefit with precipitation was larger and more consistently statistically significant than it was for the rotations where the precrop was a cereal. These findings are generally robust to sensitivity analyses where we include fixed effects to control for geographical subregions when estimating the heterogeneity of the rotation benefits with weather (Appendix A.4; Figure A5), although for cereal precrops our sensitivity analyses did not clearly suggest that the rotation benefit was higher at higher precipitation.
The direction of how the rotation benefit varies with temperature was not consistently positive or negative; however, some notable patterns exist. When soybean is the outcome crop (red bars in Figure 4), the benefit of rotation is generally greater at higher growing season temperatures. In addition, when there is a legume precrop, the benefit of rotation is generally smaller at higher growing season temperatures (top right panel in Figure 4). The only exceptions are cases where the heterogeneity coefficient is not statistically significant. Moreover, sensitivity analyses (Appendix A.4, Figure A6) demonstrate that these findings are robust to the addition of various geographical fixed effects in our model used to assess the heterogeneity of the rotation benefits.
@@ 2.3 Estimated impacts of diversified crop rotations
To estimate the impact of diverse crop rotations compared to simple rotations, we looked at the three year rotation history. Within each country (except China), and for each combination of precrop A and outcome crop B, samples where a crop other than A or B was grown before A were considered to have a diversified rotation and were deemed as treated units whereas samples where BAB was the three year rotation history were considered as the simple rotation control. Using the same causal machine learning algorithm that was used to assess 2-year rotation benefits (see Sections 4.4 and 4.6 for details), we find that diversification beyond a simple rotation confers large additional benefits in some cases, but more often the benefits are small, not statistically significant, or even negative (Table A8 (Column 7); Figure 5). Spring wheat appears to benefit most consistently from more diverse rotations, whereas when corn is the outcome crop, the estimated impact of diversification is sometimes negative, particularly for common precrop choices (soybean in Canada and the US, and winter wheat in France).
@@ 3.1 Discussion of rotation benefit estimate findings
Our results are consistent with the well established principle that switching from no rotation to crop rotation generally benefits the subsequent crop’s yields, providing some validation to our approach. The precrop effects we estimated were generally positive and significant, with two exceptions when corn was the outcome crop and soybean or wheat was the precrop, two exceptions when pasture was the precrop, and a notable exception of the SoybeanWinter Wheat sequence in the US.
@@
The negative effects for corn can be partly explained by reduced fertilizer use in rotations. For example, in the US, farmers growing SoybeanCorn are advised to use less fertilizer than those growing CornCorn [23]. Prior work [12] found that estimates of the impact of the SoybeanCorn rotation using a similar observational approach that omitted fertilizer use underestimated the rotation benefit estimates based on randomized field experiments (in the experiments, fertilizer use was typically the same in rotated and non-rotated subplots). Negative effects in Canada for cases where pasture was a precrop could be explained by the fact that PastureSoybean and PastureCorn samples often reflected a conversion of pastoral land to cropland rather than a true rotation, and thus may correspond to less productive land. Indeed, among our samples from Canada that had either a PastureCorn or a PastureSoybean sequence, roughly 64% observed a switch from a pastoral classification to a nonpastoral classification only once between 2011-2020, and 29% observed such a switch twice. Further, when using a different satellite-based proxy for yield in our sensitivity analysis in Appendix A.2, the negative estimated effects of the PastureCorn and PastureSoybean rotations were no longer statistically significant.
@@
We estimated a large negative effect of the SoybeanWinter Wheat rotation on subsequent winter wheat yields in the United States. One possible explanation for this is that winter wheat has higher yields at earlier planting dates [24], but the SoybeanWinter Wheat rotation can prevent early planting of winter wheat. According to a USDA report [25], the typical winter wheat planting window started and ended earlier than the typical soybean harvest window in most of the states that were highly represented in our study’s SoybeanWinter Wheat and Winter WheatWinter Wheat samples (Figure A10). In addition, some experiments have suggested that soybean precrops lead to reduced soil water content in the subsequent spring because their residues are ineffective at trapping snow compared to the residues of cereals such as corn and sorghum [26]. Meanwhile, residues from winter wheat (typically harvested in the summer) are effective at trapping snow (in the subsequent winter), particularly when the residues are left standing [27]. Therefore, in semi-arid regions with snow and in fields that use favorable residue management practices, a winter wheat precrop leads to more soil water availability in the subsequent spring than a soybean precrop does. Low yields in the SoybeanWinter Wheat rotation have also been reported in the experimental literature, although the results are mixed. A field experiment in the north China plain found that the SoybeanWinter Wheat rotation led to lower Winter Wheat yields than in a CornWinter Wheat rotation, particularly in dry years [28] which the authors attributed to soybean reducing the soil available water for subsequent wheat growth. Contrary to our findings, an experiment in Woodslee, Ontario found higher winter wheat yields in a Soybean-Winter Wheat rotation than in monocropped winter wheat [29]; however, the study was conducted in a location with high mean annual precipitation, and to our knowledge, the treated and control plots had identical planting dates. The fact that some SoybeanWinter Wheat versus Winter WheatWinter Wheat experiments use the same planting dates for the treatment and control subplots and are conducted in highly favorable weather conditions for the rotation serves as an example of how our observational approach can be used to uncover external validity issues in the experimental crop rotation literature.
@@
We also find that some rotations are particularly beneficial for the outcome crop. For example, legumes appear particularly effective at improving subsequent spring wheat yields in the US and Canada (Figure A1, Table A8). This is consistent with field experiments in the region, which have shown that a variety of legumes, including soybeans and dry beans, increase nitrogen availability and yield of the subsequent spring wheat crop [30, 31]. A global metanalysis of randomized field experiments found legumes to be the most effective precrops for wheat [32], although the study did not distinguish between spring and winter wheat.

### v2

# Precrop Payoffs: Causal machine learning reveals large but variable yield benefits of crop rotation in major breadbaskets

@@ Abstract
Building sustainable food systems that are resilient to climate change will require improved agricultural management and policy. One common practice that is well-known to benefit crop yields is crop rotation, yet there remains limited understanding of how the benefits of crop rotation vary for different crop sequences and for different weather conditions. To address these gaps, we leverage crop type maps, satellite data, and causal machine learning to study how precrop effects on subsequent yields vary with cropping sequence choice and weather. Complementing and going beyond what is known from randomized field trials, we find that (i) for those farmers who do rotate, the most common precrop choices tend to be among the most beneficial, (ii) the effects of switching from a simple rotation (which alternates between two crops) to a more diverse rotation were typically small and sometimes even negative, (iii) precrop effects tended to be greater under rainier conditions, (iv) precrop effects were greater under warmer conditions for soybean yields but not for other crops, and (v) legume precrops conferred smaller benefits under warmer conditions. Our results and the methods we use can enable farmers and policy makers to identify which rotations will be most effective at improving crop yields in a changing climate.
@@ 1 Introduction
Crop rotation is a well-established agronomic practice that typically raises crop yields while reducing the need for fertilizer and pesticide inputs. Yield benefits can arise from several mechanisms, including suppression of harmful weeds, pests, and diseases [1, 2], increase of soil availability of essential nutrients such as nitrogen [3, 4, 5], and improved soil health and microbial biomass [6, 7]. Although these mechanisms are well established, less is known about how the benefits of crop rotation depend on the specific cropping sequence and on the weather conditions. This knowledge gap makes it difficult to assess how sub-optimal current farming practices are (in many regions farmers often grow the same crop each year or use a simple 2-crop rotation) and whether shifts in crop rotation could help farmers adapt to ongoing climate changes. Answering these questions is crucial but will require moving beyond traditional research approaches that leverage data from multiple randomized field experiments.
@@
Recent advances in crop type mapping, satellite imagery, and causal machine learning have enabled researchers to assess the benefit of crop rotations using much larger sample sizes which allow examination of how rotation benefits vary with weather, rotational diversity, and the specific precrop and outcome crop. A number of recent studies used crop type maps and satellite imagery to assess the benefit of crop rotations [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and the accuracy of some of these observational approaches have been validated against data from randomized field experiments [12] or expert recommendations [18]. However, all of these studies have limitations in terms of their scope, their methodology, or both (see Table S6 for details). In summary, none of these studies assess the impact of switching from a simple rotation to a diverse rotation, and only a few of them assess how the benefit of rotation varies with weather [12, 19, 17, 20]. Moreover, most of these studies do not use a formal causal inference framework, and instead often rely on simple comparisons between rotated and non-rotated samples. Only [12, 19, 20] use a formal causal inference framework by leveraging modern causal machine learning tools. Of these studies, only [20] quantifies the uncertainty in their estimates of how the benefit of rotation varies with weather. As with the current study, most of these studies focus on precrop effects (defined as the effect of the previous year’s crop type on the yield of the current crop) rather than total or long-term yield effects of rotation on all crops in the sequence. The former is a major (but not the only) component of rotation effects, while the latter are difficult to detect given the limited number of years in which large-scale, high-resolution crop type maps are available.
@@
2 Overview of methods
This section gives an overview of the data and methods. For a more detailed description of the data and methods used and for explanations of some of the concepts and choices in the analysis, see the supplementary Materials and Methods section.
2.1 Dataset and rotations studied
We used annual crop type maps from northeastern China [24] (2017–2019), Canada [25] (2011–2020), France [26] (2015–2021), and the United States [27] (2008–2021). Using these crop maps and Google Earth Engine (GEE), we extracted a random sample of geographical points among croplands of interest. At each sampled location, monthly precipitation, monthly maximum and minimum temperature averages, and monthly vapor pressure deficit (VPD) were extracted from the TerraClimate gridded weather dataset [28]. For each sampled location and year, we extracted a time series of various spectral and quality assessment bands from Sentinel-2 (in China and France) and Landsat (in Canada and the US). For more details on the data sources, how they were harmonized across multiple spatial resolutions, and the sampling scheme, see Section A.1.
For each sampled location and year, the time series of multispectral satellite data was converted to a normalized measure of peak greenness, which serves as a proxy for crop yield (see Section A.2 for details). In summary, we first computed a time series of the Green Chlorophyll Vegetation Index (GCVI) [29]. We then estimated the peak GCVI in each location and year by removing observations that were corrupted by cloud cover (according to the satellites’ quality assessment bands), fitting a harmonic regression, and calculating the peak of the best fit curve. GCVI was originally designed to capture chlorophyll content in crops [29] and peak GCVI has been shown to be a useful predictor of crop yields [30, 31, 32, 33]. For increased interpretability, estimated peak GCVI values were normalized to have mean for each crop type in each country.
For each location and year in our sample, we defined the precrop to be the crop grown in the previous year and the outcome crop to be the crop grown in the current year. Samples where the precrop and outcome crop were the same were classified as having no rotation and were used as control units. Samples where the outcome crop B and precrop A were different were considered treated units and were categorized as having an AB cropping sequence. In each country, we restricted our attention to at most four outcome crops of interest and for each outcome crop and country we studied the impact of the most prevalent precrops in the sample (Figure 2). See Section A.3 for more details about the precrop and outcome crop selection criteria used.
2.2 Estimating precrop and diversification effects
We fit causal forests [22, 34] to estimate the precrop effects. A causal forest is a recently developed method for conducting causal inference in observational studies, and it can be used to estimate the average treatment effect as well as the treatment effect as a function of the covariates. Causal forests have been used to study the effects of various agronomic practices such as tillage [35], cover cropping [36], and crop rotation [12, 20] on crop yields as well as the impact of crop rotation on net primary productivity [19]. In the context of assessing the effect of the SoybeanCorn crop rotation on corn yield, rotation benefits estimated by fitting a causal forest to a satellite-derived dataset were found to have a statistically significant positive correlation with estimates from actual randomized field experiments [12].
For each of four countries, and for each common combination of precrop A and outcome crop B, we conducted the following procedure to estimate the impact of the AB sequence compared to the BB sequence on a proxy for the yield of B in the second year. First, we subsetted our data to only include samples with AB or BB sequences and designated a binary treatment variable to be for samples with AB sequences and for samples with BB sequences. Next, we set the outcome variable to be the normalized peak GCVI variable described in Section A.2, which serves as a proxy for crop yield. We then let denote a vector of year, latitude, longitude, 7 weather covariates, and (in the US only) irrigation status (from the Landsat-based Irrigation Dataset [37]). The weather covariates captured early precipitation, growing season precipitation, average maximum and minimum daily temperatures in the growing season, and VPD for the three peak months of the growing season. We then fit a causal forest, using the grf package in R [38], to estimate the function
In words, is the difference in the mean normalized peak GCVI in the treated group minus that in the control group written as a function of the covariates . Under the assumption that there are no unmeasured confounders, is equal to the conditional average treatment effect (which is the mean causal effect of the treatment on the outcome as a function of ). Finally the average precrop effects were estimated by taking an overlap-weighted average of the estimated values, assigning little weight to locations in space and time where only or only was observed and could not have been properly estimated.
We refer readers to Section A.4 for a more explicit description of the causal forest approach, the parameters and software used, the estimand, how standard errors were calculated, and the specific weather covariates used. That section describes various justifications and robustness checks regarding the choice of outcome variable, the choice of an overlap-weighted estimand, and our choice to not account for spatial correlations when computing standard errors (see also Appendix C.5). We also compared some of the causal forest results to those from a simpler approach of calculating within each 1°1° grid cell the difference between the mean of in the treated (rotated) versus control (non-rotated) group, finding similar qualitative results (see Appendix C.6). For model fit accuracy metrics, see Appendix K and Table S8.
To estimate the impact of diverse crop rotations compared to simple rotations, we used the same causal forest methodology with a different choice of treatment variable that was based on the three year rotation history. Within each country (except China), and for each combination of precrop A and outcome crop B, samples where a crop other than A or B was grown before A were considered to have a diversified rotation and were deemed as treated units whereas samples where BAB was the three year rotation history were considered as the simple rotation control units . See Section A.6 for more details.
2.3 Estimating heterogeneity of precrop effects with weather
To assess how the benefit of the AB sequence versus the control BB varied with weather, we ran the following linear regression of on the year , the growing season precipitation , and the growing season temperature
The above regression was implemented using the best_linear_projection function in the grf package [38], which provided standard errors and -values that appropriately accounted for the fact that the function was unknown and had to be estimated from the data. The regression was fit on the subsample of points where the two year sequence was either AB or BB and excluded regions where the estimated propensity score was between and (allowing us to ignore regions where only AB or only BB was observed, but not both). For more details about the subsample used, the growing season weather variables, how confidence intervals were constructed, and for a justification of the linear modeling choices see Section A.5. As a supplementary sensitivity analysis, variants of the above regression model that included geographical fixed effects were also considered (Appendix C.4).
To rescale the estimated regression coefficients to be on a more interpretable scale, on the same subsample, we computed the 25th percentile and 75th percentile for growing season precipitation and for growing season temperatures denoted by , , , and , respectively. Letting and denote the estimated regression coefficients from the above regression, our rescaled heterogeneity coefficients for temperature and weather were calculated with
The rescaled heterogeneity coefficients can be roughly interpreted as the amount by which the estimated benefit of rotation on our normalized yield proxy are expected to change as the growing season precipitation (or temperature) increases from the 25th percentile to the 75th percentile in the regions where both the rotation AB and the no rotation control BB are commonly observed. The above process was repeated for each precrop A, outcome crop B, and country listed in Table S9.
2.4 Calibration with crop yield data
To improve the interpretability of our results, we converted all estimated precrop effects, diversification effects, and heterogeneity coefficients from a unitless GCVI scale to the crop yield scale by using a post-hoc, linear calibration with subnational-level crop yield data. In particular, we fit linear regressions of annual county-level crop yield data from [39] on county-level averages of normalized peak GCVI. Best fit linear regression coefficients were then used to convert all results to be in units of % of the average yield for the outcome crop and country of interest. For more details about the calibration and sensitivity analyses with alternative calibration schemes see Section A.7 and Appendix J.
@@ 3.1 Estimates of impacts of 2-year crop sequences
The results (Table S9 (Column 6); Figure 3) indicate that rotations generally benefit crop yields, but there are considerable differences in these benefits for different precrop and outcome crop combinations. Of the four outcome crops considered, the rotation benefit estimates were generally largest for spring wheat and smallest for corn. For winter wheat outcome crops, soybean precrops had a negative estimated effect whereas rapeseed or fallow precrops had a positive estimated effect. For spring wheat outcome crops, legume precrops and rapeseed precrops were generally the most effective, whereas for soybean outcome crops, cereals were the most effective types of precrops (Figure S1).
The results also suggest that beneficial precrops are being exploited by farmers. In Figure 3, we see that the most common precrops have either the highest or nearly the highest benefits. There are some exceptions in which less common precrop choices are estimated to be more effective than the most common precrop choices when spring wheat or corn is the outcome crop. In addition, only relatively common precrops are included in the analysis (Section A.3), so our analysis does not rule out the possibility that some very rare precrops are more effective than commonly used precrops.
@@ 3.2 Variability of rotation benefits with weather
The results (Table S9 (Columns 8–9); Figure 4) indicate that at higher growing season precipitation there is generally a greater benefit due to rotation. In particular, with a few exceptions in France and for fallow precrops, cases where the rotation benefit is greater at lower precipitation levels did not have statistically significant heterogeneity coefficients. In addition, based on a sensitivity analysis where we use the log of estimated crop yield as the outcome variable (Appendix C.3; Figure S5, bottom left panel), we find that this result is not merely driven by the fact that crop productivity is higher in rainier conditions. Instead, we see increases in both the relative effectiveness and absolute effectiveness of crop rotation in rainier conditions. For the rotations with a legume or broadleaf precrop, the heterogeneity of the rotation benefit with precipitation was larger and more consistently statistically significant than it was for the rotations where the precrop was a cereal. These findings are generally robust to sensitivity analyses where we include fixed effects to control for geographical subregions when estimating the heterogeneity of the rotation benefits with weather (Appendix C.4; Figure S6), although for cereal precrops our sensitivity analyses did not clearly suggest that the rotation benefit was higher at higher precipitation.
The direction of how the rotation benefit varies with temperature was not consistently positive or negative; however, some notable patterns exist. When soybean is the outcome crop (red bars in Figure 4), the benefit of rotation is generally greater at higher growing season temperatures. In addition, when there is a legume precrop, the benefit of rotation is generally smaller at higher growing season temperatures (top right panel in Figure 4). The only exceptions are cases where the heterogeneity coefficient is not statistically significant. Moreover, sensitivity analyses (Appendix C.4, Figure S7) demonstrate that these findings are robust to the addition of various geographical fixed effects in our model used to assess the heterogeneity of the rotation benefits.
@@ 3.3 Estimated impacts of diversified crop rotations
The results suggest that diversification beyond a simple rotation confers large additional benefits in some cases, but more often the benefits are small, not statistically significant, or even negative (Table S9 (Column 7); Figure 5). Spring wheat appears to benefit most consistently from more diverse rotations, whereas when corn is the outcome crop, the estimated impact of diversification is sometimes negative, particularly for common precrop choices (soybean in Canada and the US, and winter wheat in France). Supplementary results also showed that the impact of diversification beyond a simple rotation tended to be larger in rainier conditions (Figure S2), although many estimates of the heterogeneity of the diversification effects were not statistically significant due to limited sample sizes in the diversification analyses.
@@ 4.1 Discussion of rotation benefit estimate findings
Our results are consistent with the well established principle that switching from no rotation to crop rotation generally benefits the subsequent crop’s yields, providing some validation to our approach. For corn yields, our results suggest that winter wheat and soybean were beneficial precrops in the United States and Canada, although this conclusion could not be drawn in China and France or for spring wheat precrops. For soybean yields, corn was a particularly beneficial precrop (in the US, Canada, and China), while wheat (in the US and Canada) and rapeseed (in Canada) precrops were found to have smaller, yet still positive benefits on yields. For spring wheat yields in the US and Canada, rapeseed and legumes such as soybeans, peas, lentils and drybeans were found to be particularly beneficial precrops, whereas in the US sunflower and corn precrops had more modest benefits. For winter wheat yields in France, rapeseed was again found to be a particularly effective precrop, and sunflower, and chard/beet, corn silage, and corn, were also quite effective as precrops, with the latter finding for corn precrops also holding in the US. Overall, the precrop effects that we estimated were generally positive and statistically significant, with two exceptions when corn was the outcome crop and soybean or wheat was the precrop, two exceptions when pasture was the precrop, and a notable exception of the SoybeanWinter Wheat sequence in the US.
@@
The negative effects for corn can be partly explained by reduced fertilizer use in rotations. For example, in the US, farmers growing SoybeanCorn are advised to use less fertilizer than those growing CornCorn [40]. Prior work [12] found that estimates of the impact of the SoybeanCorn rotation using a similar observational approach that omitted fertilizer use underestimated the rotation benefit estimates based on randomized field experiments (in the experiments, fertilizer use was typically the same in rotated and non-rotated subplots). Omission of fertilizer also explains why the estimated precrop effects of soybean on corn were much smaller than the estimated precrop effects of corn on soybean, contrary to findings from randomized field experiments. Negative effects in Canada for cases where pasture was a precrop could be explained by the fact that PastureSoybean and PastureCorn samples often reflected a conversion of pastoral land to cropland rather than a true rotation, and thus may correspond to less productive land. Indeed, among our samples from Canada that had either a PastureCorn or a PastureSoybean sequence, roughly 64% observed a switch from a pastoral classification to a nonpastoral classification only once between 2011-2020, and 29% observed such a switch twice. Further, when using a different satellite-based proxy for yield in our sensitivity analysis in Appendix C.2, the negative estimated effects of the PastureCorn and PastureSoybean rotations were no longer statistically significant.
@@
We estimated a large negative effect of the SoybeanWinter Wheat rotation on subsequent winter wheat yields in the United States. One possible explanation for this is that winter wheat has higher yields at earlier planting dates [41], but the SoybeanWinter Wheat rotation can prevent early planting of winter wheat. According to a USDA report [42], the typical winter wheat planting window started and ended earlier than the typical soybean harvest window in most of the states that were highly represented in our study’s SoybeanWinter Wheat and Winter WheatWinter Wheat samples (Figure S12). In addition, some experiments have suggested that soybean precrops lead to reduced soil water content in the subsequent spring because their residues are ineffective at trapping snow compared to the residues of cereals such as corn and sorghum [43]. Meanwhile, residues from winter wheat (typically harvested in the summer) are effective at trapping snow (in the subsequent winter), particularly when the residues are left standing [44]. Therefore, in semi-arid regions with snow and in fields that use favorable residue management practices, a winter wheat precrop leads to more soil water availability in the subsequent spring than a soybean precrop does. Low yields in the SoybeanWinter Wheat rotation have also been reported in the experimental literature, although the results are mixed. A field experiment in the north China plain found that the SoybeanWinter Wheat rotation led to lower Winter Wheat yields than in a CornWinter Wheat rotation, particularly in dry years [45] which the authors attributed to soybean reducing the soil available water for subsequent wheat growth. Contrary to our findings, an experiment in Woodslee, Ontario found higher winter wheat yields in a Soybean-Winter Wheat rotation than in monocropped winter wheat [46]; however, the study was conducted in a location with high mean annual precipitation, and to our knowledge, the treated and control plots had identical planting dates. The fact that some SoybeanWinter Wheat versus Winter WheatWinter Wheat experiments use the same planting dates for the treatment and control subplots and are conducted in highly favorable weather conditions for the rotation serves as an example of how our observational approach can be used to uncover external validity issues in the experimental crop rotation literature.
@@
We also find that some rotations are particularly beneficial for the outcome crop. For example, legumes appear particularly effective at improving subsequent spring wheat yields in the US and Canada (Figure S1, Table S9). This is consistent with field experiments in the region, which have shown that a variety of legumes, including soybeans and dry beans, increase nitrogen availability and yield of the subsequent spring wheat crop [47, 48]. A global metanalysis of randomized field experiments found legumes to be the most effective precrops for wheat [49], although the study did not distinguish between spring and winter wheat.